# Physics-informed machine learning for combustion: A review

### v1

@@ 4.3.2 Kinetics discovery
In terms of embedding the knowledge of chemical kinetics into NNs, the chemical reaction neural network (CRNN), proposed by Ji et al. 75 for kinetics discovery, is an earlier work than References 117, 118, 119 and is also trained in a NODE-way. As shown in Figure 15, a neuron of CRNN represents a single-step reaction, where the law of mass action (Equation (3.28)) and Arrhenius law (Equation (3.32)) are embedded as the nodes and weights. For the input layer, the weights represent the reaction orders and the bias represents the rate constant in the logarithmic scale. The weights of the output layer correspond to the stoichiometric coefficients. A multi-step reaction system can be represented by stacking the neurons to formulate a CRNN with one hidden layer, where the number of hidden neurons is equal to the number of reactions. Due to the physical interpretability, once the CRNN has been trained, the kinetic parameters can be easily recovered by checking the weights and biases, then the reaction pathways can be fully revealed. The number of hidden nodes/reactions, which is a hyperparameter, can be determined by grid searching and a threshold pruning strategy is suggested to clip the NN weights below a certain threshold 75. The CRNN was first tested on three cases with increasing complexity: an elementary reaction network without temperature dependence, biodiesel production with temperature dependence, and an enzyme reaction network in which some species are present in both reactants and products. Results demonstrated that CRNN could accurately learn the reaction pathways and kinetic parameters from noisy data and even from incomplete data with missing species.
Figure 15: Schematic of the CRNN illustrated for a reaction system with four species and four reaction steps. Reprinted from Reference 75, with permission.
Later, CRNN has been applied to many reaction systems, such as the biomass pyrolysis 293 and the decomposition of energetic materials 294, 295, 296, 297. Li et al. 298 developed the Bayesian CRNN to simultaneously infer the reaction pathways and perform uncertainty quantification on the identified reaction pathways, kinetic parameters, and concentration predictions with unseen initial conditions. Su et al. 281 used NODEs for kinetic parameter optimization of hydrocarbon fuels, where the framework can be viewed as CRNN with only kinetic parameters unknown, all pathways fixed and all other unnecessary connections trimmed. Considering that CRNN does not ensure mass and element/atom conservations, Doeppel et al. 282 developed the atom conserving CRNN (AC-CRNN), where the atom conservation was enforced by adding a NN layer that multiplicated the key species coefficients with a conservation matrix. The standard cases in Reference 75 was tested. Compared to CRNN, the AC-CRNN exhibited improved stability and robustness against noisy and missing data, and was more efficient due to the reduced number of trainable parameters.
@@ 5 PIML for combustion reacting flows
Choi 299, 2022.1 3D+t Inverse Mass, $\mathbf{U}$, $T$, $Y_{1-3}$ No Van de Vusse Sim. MLP
Wang 260, 2022.7 2D Forward $Z$ No Absent (fast chemistry) Sim. MLP
Sitte 300, 2022.8 2D+t Inverse Mass, $\mathbf{U}$ No Absent Sim. MLP
Liu 301, 2023.9 1D, 2D, 2D+5 Forward Mass, $\mathbf{U}$, $T$, $Y_{1-5}$ No 1-step CH4 Sim. MLP
Song 302, 2024.10 2D Both Mass, $\mathbf{U}$, $Z$, $c$ No Detailed CH4 (FPV table) Sim. MLP
Cao 303, 2024.11 2D Forward Mass, $\mathbf{U}$, $T$, $Y_{1-5}$ No 1-step CH4 Sim. MLP
Wu 112, 2025.1 1D (1) Both (2) Forward (3) Inverse (1) $T$ (2) $T$, $Y_{1-5}$ (3) Mass, $V$, $T$, $Y_{1-5}$ No (1) Simplified FRC (2)(3) 1-step CH4 Sim. (1) MLP (2)(3) MMLP-RWF URL
von Saldern 226, 2022.11 2D Inverse Mass Filtered Absent Sim. MLP
Liu 304, 2023.12 3D+t Inverse Mass, $\mathbf{U}$, $T$, $_{1-6}, Y\sum Y$ Direct 2-step CH4 Sim. MLP
Peng 305, 2024.4 2D+t Inverse Mass, $\widetilde{\mathbf{U}}$, $\widetilde{T}$, $k$, $\varepsilon$ Mean Absent Sim. MLP
Taassob 306, 2023.10 2D+1 Inverse Mass, $\widetilde{\mathbf{U}}$, $\widetilde{\mathrm{PC}}_{1-2}$, $\widetilde{Z}$, $\mu_{t}$ Mean (learned) Learned Exp. MLP
Taassob 106, 2024.6 2D+2 Inverse Mass, $\widetilde{\mathbf{U}}$, $\widetilde{Y}_{1-6}$, $\widetilde{Z}$, $\mu_{t}$ Mean (learned) Learned Exp. DeepONet
Liu 107, 2024.6 (1) 2D+t (2) 2D Inverse (1) $\mathbf{U}$, $T$ (2) $Y_{\mathrm{soot}}$ No (1) Absent (2) Learned Exp. (1) MLP (2) AE-DeepONet
Yadav 307, 2024.12 (1) 1D (2)(3) 2D Inverse Mass, (1) $c$ ($\mathbf{U}$) (2) $c$ (3) $\widetilde{c}$ (1)(2) No (3) Mean (learned) (1)(2) Pre-learned (3) EBU model (1)(2) Sim. (3) Exp. MLP
Wang 308, 2023.7 1D+t Inverse Mass, $u$, $E$, $c$ No Simplified FRC Sim. MLP
Deng 309, 2024.7 2D+t+2 Inverse Mass, $\widetilde{\mathbf{U}}$ Mean Absent Sim. ResUNet
Bode 310, 2021.1 3D Inverse Mass Filtered to full Absent Sim. ESR-GAN URL
Bode 311, 2023.2 3D+t Inverse Mass, $\sum Y$ Filtered to full Absent Sim. ESR-GAN
Wang 312, 2023.11 2D Inverse Mass, $\mathbf{U}$ No Absent Sim. GAN
Wang 313, 2024.8 2D+t Inverse Mass No Absent Sim. R-GAN URL
@@
which describes the process of phase separation in multi-component alloy systems, and has become a benchmark problem for validating PINN algorithms 66, 173, 175, 174. It can be seen that the reaction source term is quite different from the combustion ones (Section 3.4), and only one variable is considered. Other DR systems, such as $\lambda$-$\omega$, FitzHugh-Nagumo, and Gray-Scott equations have been solved by PhyCRNet 121 and PeRCNN 73. Although these problems involve more than one variable, the polynomial reaction kinetics still differ from the chemical kinetics of practical combustion. Nevertheless, these studies have validated the feasibility of PINNs to solve PDEs with both diffusion terms and highly nonlinear reaction source terms. Niaki et al. 314 used PINNs to solve the DR PDEs of the thermochemical curing process of composite-tool systems during manufacture, where the cure kinetics contained the Arrhenius law, further validating the applicability of PINNs to solve combustion-like DR problems.
@@
For ADR problems, their difference from CDR problems is that the velocities are constants or functions instead of field variables, so the nonlinearity is mainly caused by reaction instead of convection. Ngo et al. 315 applied PINNs to the solution and parameter inference of a fixed-bed reactor model for catalytic CO2 methanation. The momentum and energy equations were neglected due to the low pressure drop and isothermal conditions, respectively, so the governing equations were 1D advection-reaction equations of four species, which considered the law of mass action and Arrhenius law. Results showed that the forward PINN could solve the reactor model, whereas the inverse PINN could reveal an unknown effectiveness factor using the pre-trained forward PINN. Hou et al. 316 applied PINNs to forward and inverse one-dimensional one-component (1D1C) ADR problems, where both location-dependent and linear reaction terms were considered. The adopted first derivative constraint (FDC) method, which is similar to gPINN 176, has been demonstrated to enhance the model stability and accuracy. To avoid the slow training speed caused by FDC, Sun et al. 317 proposed an NN connection structure named deep function family construction, which added residual connections to fully connected NNs. The modified PINN was applied to forward and inverse 1D3C ADR problems, where the reaction source terms were given by the law of mass action for a reaction system of three species and five reactions. Both steady and unsteady problems were solved and the modified PINN outperformed the vanilla PINN, especially for cases with steep fronts. The results of inverse problems indicated the capability of PINNs to learn missing physics (multiple unknown parameters) in ADR systems. This was further demonstrated by Hou et al. 318, where the inverse problem of the 1D3C ADR system was also solved. Laghi et al. 319 compared four PINNs on six 1D steady-state ADR problems with analytical solutions. The four PINNs were vanilla PINNs, PI-ELM 147, Deep-TFC 132, and X-TFC 133. Comparative analysis demonstrated that X-TFC was the best-performing framework in terms of both accuracy and computational time.
@@ 5.2 Laminar combustion
This subsection will present a chronological overview of studies on PINNs for laminar subsonic combustion. Choi et al. 299 used PINNs to reconstruct the 3D transient fields of a continuous stirred tank reactor (CSTR) with van de Vusse reaction. The governing equations of two reference frames were considered to enable the reconstruction of the multi-reference frame system. The fields can be accurately reconstructed from the data of velocity, temperature, and species, and some proposed strategies, such as the similarity-based sampling, have been shown to improve the model performance. Wang et al. 260 used PINNs to forward solve the $Z$-equation for a jet diffusion flame and obtained accurate results. Because the assumptions of unity Lewis number and infinite fast chemistry are adopted, the $Z$-equation is free of reaction source terms and the temperature and species can be determined solely by $Z$. Sitte et al. 300 used PINNs to reconstruct the velocity fields in puffing pool fires based on the observations of density, pressure, and temperature, following the HFM framework 244, as illustrated in Figure 16. The continuity and momentum equations are taken as the physical losses and a two-stage training strategy is adopted. The results show that PINNs can provide satisfactory velocity results and can act as a physics-based denoiser to smooth out the noise on the measured data.
@@
Figure 16: Applying the HFM framework to reconstruct puffing pool fires. Reprinted from Reference 300.
@@
Liu et al. 301 applied PINNs to solve forward problems of 1D and 2D premixed combustion with temperature-dependent properties and a one-step global mechanism of CH4. To the best of our knowledge, although some simplifications are adopted for the transport models and the reaction mechanism is simple, this is the first study on PINNs for combustion reacting flows with detailed physical models and chemical kinetics based on the laws of mass action and Arrhenius. More importantly, after validating the model accuracy on 1D and 2D problems, a surrogate model was constructed for the 2D combustion, where five parameters were added to the NN inputs so that the model can learn multi-condition solutions, as shown in Figure 17. The parameterized PINN surrogate model was shown to accurately predict the fields, be applicable to sensitive analysis, and be much more efficient for industrial designs than traditional numerical simulations. Song et al. 302 applied PINNs to 2D laminar premixed flames considering detailed chemistry of CH4, which was accomplished by integrating the flamelet/progress variable (FPV) model. The equations of mass, momentum, $Z$, and $c$ were solved by PINN, during which the thermophysical properties and source term of $c$ were extracted from the pre-tabulated FPV library, as illustrated in Figure 18. The model can accurately solve the fields with and without observation data, demonstrating the feasibility of PINNs to integrate detailed reaction mechanisms using tabulation methods. Using pseudo-time stepping PINNs (PTS-PINNs) 212, Cao et al. 303 solved the premixed and non-premixed flames with a one-step mechanism of CH4 and obtained satisfactory results. Vanilla PINNs failed to solve the premixed combustion with large gradients and tended to predict non-reacting solutions, but PTS-PINNs showed great convergence performance not only for the premixed combustion but also for non-premixed combustion where the thin reaction zones introduced a higher level of complexity.
@@
Figure 17: Schematic of the parameterized combustion field surrogate model based on PINNs. Reprinted from Reference 301, with permission.
@@
Figure 18: Schematic of the PINNs coupled with the FPV model. Reprinted from Reference 302, with permission.
@@
The above studies only involve forward problems or inverse problems of field reconstruction without parameter inference, which is another important aspect of inverse problems. In computational 1D laminar flames, which are important for the study of flame limit, stretch, laminar flame speed ($s_{L}$), and flamelet tabulation, unknown eigenvalues exist even in forward problems, such as $s_{L}$ in freely-propagating premixed flames 4, 7, 8. Wu et al. 112 established the FlamePINN-1D framework to solve both forward and inverse problems of 1D laminar flames in a unified manner, as illustrated in Figure 19. Three cases were solved: freely-propagating premixed flames with simplified and detailed physical models, and counterflow premixed flames with detailed physical models. The detailed physical models are the same as those in Section (3.1) to (3.5) with a one-step global mechanism of CH4 being considered, except that the governing equations are for 1D reacting stagnation flows. A warmup pretraining strategy and some hard physical constraints were proposed. For example, the predicted species mass fractions are transformed by the Softmax function:
@@
where $Y_{k}^{\#}$ is the output of NN. The transformation ensures that the range constraint ($0\leqslant Y_{k}\leqslant 1$) and species conservation ($\sum_{k}Y_{k}=1$, Equation (3.1)) can be exactly satisfied, so it can be used for various combustion problems. $Y^{\#}$ can be viewed as the logarithmic normalization of $Y$ 268, which can mitigate the multiscale of their magnitudes. The results show that for forward problems, FlamePINN-1D can accurately solve the flame fields and simultaneously infer the $s_{L}$ of freely-propagating premixed flames. For inverse problems, FlamePINN-1D can accurately reconstruct the continuous fields and infer the transport properties and chemical kinetics parameters from noisy sparse observations, which indicates the possibility of optimizing chemical mechanisms from observations of laboratory 1D flames. In fact, in their inverse problems where the activation energy was unknown, the chemical kinetics part can be considered as a CRNN with other parameters are known. It is proven that the model tends to get unburned solutions without warmup pretraining, the similar situation to Reference 303. Other adopted strategies, such as hard constraints, thin-layer normalization 85, and modified MLP with random weight factorization (MMLP-RWF) 79, have also been proven to be essential for the model performance.
Figure 19: FlamePINN-1D for solving forward and inverse problems of 1D laminar flames. (a) Schematic of the FlamePINN-1D framework. (b) Case 1: freely-propagating premixed flames with simplified physical models. (c) Case 2: freely-propagating premixed flames with detailed physical models. (d) Case 3: counterflow premixed flames with detailed physical models. Reprinted from Reference 112, with permission.
In addition to the classical PINNs, the PINNs in a broad sense (Table 1) can also be applied to laminar combustion problems. For example, Corrochano et al. 320 developed a physics-aware ROM for an axisymmetric laminar CH4-O2 flame. The simulation data were preprocessed to obtain the proper orthogonal decomposition (POD) modes, then deep learning models were used as a time integrator to advance temporal modes in time, i.e., the models took the AR paradigm instead of the NAR paradigm adopted in classical PINNs (see Section 2.3.1). A physics-aware loss function was introduced which constrained the predicted species mass to be equal to the true ones in each time instant. This loss differed from the classical PINN losses in that it was still a data-driven loss without PDE information, so it was called "physics-aware" rather than "physics-informed", but can still be called "physics-informed" in a broad sense in our extended definition. Two models, including the LSTM and 1D CNN, were tested, demonstrating that the physics-aware loss could slightly improve the model performance.

### v2

@@ 4.3.2 Kinetics discovery
In terms of embedding the knowledge of chemical kinetics into NNs, the chemical reaction neural network (CRNN), proposed by Ji et al. 78 for kinetics discovery, is an earlier work than References 120, 121, 122 and is also trained in a NODE-way. As shown in Figure 15, a neuron of CRNN represents a single-step reaction, where the law of mass action (Equation (3.28)) and Arrhenius law (Equation (3.32)) are embedded as the nodes and weights. For the input layer, the weights represent the reaction orders and the bias represents the rate constant in the logarithmic scale. The weights of the output layer correspond to the stoichiometric coefficients. A multi-step reaction system can be represented by stacking the neurons to formulate a CRNN with one hidden layer, where the number of hidden neurons is equal to the number of reactions. Due to the physical interpretability, once the CRNN has been trained, the kinetic parameters can be easily recovered by checking the weights and biases, then the reaction pathways can be fully revealed. The number of hidden nodes/reactions, which is a hyperparameter, can be determined by grid searching and a threshold pruning strategy is suggested to clip the NN weights below a certain threshold 78. The CRNN was first tested on three cases with increasing complexity: an elementary reaction network without temperature dependence, biodiesel production with temperature dependence, and an enzyme reaction network in which some species are present in both reactants and products. Results demonstrated that CRNN could accurately learn the reaction pathways and kinetic parameters from noisy data and even from incomplete data with missing species.
Figure 15: Schematic of the CRNN illustrated for a reaction system with four species and four reaction steps. Reprinted from Reference 78, with permission.
Later, CRNN has been applied to many reaction systems, such as the biomass pyrolysis 296 and the decomposition of energetic materials 297, 298, 299, 300. Li et al. 301 developed the Bayesian CRNN to simultaneously infer the reaction pathways and perform uncertainty quantification on the identified reaction pathways, kinetic parameters, and concentration predictions with unseen initial conditions. Su et al. 284 used NODEs for kinetic parameter optimization of hydrocarbon fuels, where the framework can be viewed as CRNN with only kinetic parameters unknown, all pathways fixed and all other unnecessary connections trimmed. Considering that CRNN does not ensure mass and element/atom conservations, Doeppel et al. 285 developed the atom conserving CRNN (AC-CRNN), where the atom conservation was enforced by adding a NN layer that multiplicated the key species coefficients with a conservation matrix. The standard cases in Reference 78 was tested. Compared to CRNN, the AC-CRNN exhibited improved stability and robustness against noisy and missing data, and was more efficient due to the reduced number of trainable parameters.
@@ 5 PIML for combustion reacting flows
Choi 305, 2022.1 3D+t Inverse Mass, $\mathbf{U}$, $T$, $Y_{1-3}$ No Van de Vusse Sim. MLP
Wang 263, 2022.7 2D Forward $Z$ No Absent (fast chemistry) Sim. MLP
Sitte 306, 2022.8 2D+t Inverse Mass, $\mathbf{U}$ No Absent Sim. MLP
Liu 307, 2023.9 1D, 2D, 2D+5 Forward Mass, $\mathbf{U}$, $T$, $Y_{1-5}$ No 1-step CH4 Sim. MLP
Song 308, 2024.10 2D Both Mass, $\mathbf{U}$, $Z$, $c$ No Detailed CH4 (FPV table) Sim. MLP
Cao 309, 2024.11 2D Forward Mass, $\mathbf{U}$, $T$, $Y_{1-5}$ No 1-step CH4 Sim. MLP
Wu 115, 2025.1 1D (1) Both (2) Forward (3) Inverse (1) $T$ (2) $T$, $Y_{1-5}$ (3) Mass, $V$, $T$, $Y_{1-5}$ No (1) Simplified FRC (2)(3) 1-step CH4 Sim. (1) MLP (2)(3) MMLP-RWF URL
von Saldern 229, 2022.11 2D Inverse Mass Filtered Absent Sim. MLP
Peng 311, 2024.4 2D+t Inverse Mass, $\widetilde{\mathbf{U}}$, $\widetilde{T}$, $k$, $\varepsilon$ Mean Absent Sim. MLP
Liu 310, 2023.12 3D+t Inverse Mass, $\mathbf{U}$, $T$, $_{1-6}, Y\sum Y$ Direct 2-step CH4 Sim. MLP
Liu 304, 2025.9 3D+t Inverse $\mathbf{U}$ Direct Absent Sim. MLP
Taassob 312, 2023.10 2D+1 Inverse Mass, $\widetilde{\mathbf{U}}$, $\widetilde{\mathrm{PC}}_{1-2}$, $\widetilde{Z}$, $\mu_{t}$ Mean (learned) Learned Exp. MLP
Taassob 109, 2024.6 2D+2 Inverse Mass, $\widetilde{\mathbf{U}}$, $\widetilde{Y}_{1-6}$, $\widetilde{Z}$, $\mu_{t}$ Mean (learned) Learned Exp. DeepONet
Liu 110, 2024.6 (1) 2D+t (2) 2D Inverse (1) $\mathbf{U}$, $T$ (2) $Y_{\mathrm{soot}}$ No (1) Absent (2) Learned Exp. (1) MLP (2) AE-DeepONet
Yadav 313, 2024.12 (1) 1D (2)(3) 2D Inverse Mass, (1) $c$ ($\mathbf{U}$) (2) $c$ (3) $\widetilde{c}$ (1)(2) No (3) Mean (learned) (1)(2) Pre-learned (3) EBU model (1)(2) Sim. (3) Exp. MLP
Zhang 303, 2025.9 2D(+t) Inverse Mass, $\mathbf{U}$, $T$, $_{1-5}, Y\sum Y$ (1) No (2) Mean (learned) 1-step CH4 and learned (1) Sim. (2) Exp. MRF-CNN
Wang 314, 2023.7 1D+t Inverse Mass, $u$, $E$, $c$ No Simplified FRC Sim. MLP
Deng 315, 2024.7 2D+t+2 Inverse Mass, $\widetilde{\mathbf{U}}$ Mean Absent Sim. ResUNet
Bode 316, 2021.1 3D Inverse Mass Filtered to full Absent Sim. ESR-GAN URL
Bode 317, 2023.2 3D+t Inverse Mass, $\sum Y$ Filtered to full Absent Sim. ESR-GAN
Wang 318, 2023.11 2D Inverse Mass, $\mathbf{U}$ No Absent Sim. GAN
Wang 319, 2024.8 2D+t Inverse Mass No Absent Sim. R-GAN URL
@@
which describes the process of phase separation in multicomponent alloy systems, and has become a benchmark problem for validating PINN algorithms 69, 176, 178, 177. It can be seen that the reaction source term is quite different from the combustion ones (Section 3.4), and only one variable is considered. Other DR systems, such as $\lambda$-$\omega$, FitzHugh-Nagumo, and Gray-Scott equations have been solved by PhyCRNet 124 and PeRCNN 76. Although these problems involve more than one variable, the polynomial reaction kinetics still differ from the chemical kinetics of practical combustion. Nevertheless, these studies have validated the feasibility of PINNs to solve PDEs with both diffusion terms and highly nonlinear reaction source terms. Niaki et al. 320 used PINNs to solve the DR PDEs of the thermochemical curing process of composite-tool systems during manufacture, where the cure kinetics contained the Arrhenius law, further validating the applicability of PINNs to solve combustion-like DR problems.
@@
For ADR problems, their difference from CDR problems is that the velocities are constants or functions instead of field variables, so the nonlinearity is mainly caused by reaction instead of convection. Ngo et al. 321 applied PINNs to the solution and parameter inference of a fixed-bed reactor model for catalytic CO2 methanation. The momentum and energy equations were neglected due to the low pressure drop and isothermal conditions, respectively, so the governing equations were 1D advection-reaction equations of four species, which considered the law of mass action and Arrhenius law. Results showed that the forward PINN could solve the reactor model, whereas the inverse PINN could reveal an unknown effectiveness factor using the pre-trained forward PINN. Hou et al. 322 applied PINNs to forward and inverse one-dimensional one-component (1D1C) ADR problems, where both location-dependent and linear reaction terms were considered. The adopted first derivative constraint (FDC) method, which is similar to gPINN 179, has been demonstrated to enhance the model stability and accuracy. To avoid the slow training speed caused by FDC, Sun et al. 323 proposed an NN connection structure named deep function family construction, which added residual connections to fully connected NNs. The modified PINN was applied to forward and inverse 1D3C ADR problems, where the reaction source terms were given by the law of mass action for a reaction system of three species and five reactions. Both steady and unsteady problems were solved and the modified PINN outperformed the vanilla PINN, especially for cases with steep fronts. The results of inverse problems indicated the capability of PINNs to learn missing physics (multiple unknown parameters) in ADR systems. This was further demonstrated by Hou et al. 324, where the inverse problem of the 1D3C ADR system was also solved. Laghi et al. 325 compared four PINNs on six 1D steady-state ADR problems with analytical solutions. The four PINNs were vanilla PINNs, PI-ELM 150, Deep-TFC 135, and X-TFC 136. Comparative analysis demonstrated that X-TFC was the best-performing framework in terms of both accuracy and computational time.
@@ 5.2 Laminar combustion
This subsection will present a chronological overview of studies on PINNs for laminar subsonic combustion. Choi et al. 305 used PINNs to reconstruct the 3D transient fields of a continuous stirred tank reactor (CSTR) with van de Vusse reaction. The governing equations of two reference frames were considered to enable the reconstruction of the multi-reference frame system. The fields can be accurately reconstructed from the data of velocity, temperature, and species, and some proposed strategies, such as the similarity-based sampling, have been shown to improve the model performance. Wang et al. 263 used PINNs to forward solve the $Z$-equation for a jet diffusion flame and obtained accurate results. Because the assumptions of unity Lewis number and infinite fast chemistry are adopted, the $Z$-equation is free of reaction source terms and the temperature and species can be determined solely by $Z$. Sitte et al. 306 used PINNs to reconstruct the velocity fields in puffing pool fires based on the observations of density, pressure, and temperature, following the HFM framework 247, as illustrated in Figure 16. The continuity and momentum equations are taken as the physical losses and a two-stage training strategy is adopted. The results show that PINNs can provide satisfactory velocity results and can act as a physics-based denoiser to smooth out the noise on the measured data.
@@
Figure 16: Applying the HFM framework to reconstruct puffing pool fires. Reprinted from Reference 306.
@@
Liu et al. 307 applied PINNs to solve forward problems of 1D and 2D premixed combustion with temperature-dependent properties and a one-step global mechanism of CH4. To the best of our knowledge, although some simplifications are adopted for the transport models and the reaction mechanism is simple, this is the first study on PINNs for combustion reacting flows with detailed physical models and chemical kinetics based on the laws of mass action and Arrhenius. More importantly, after validating the model accuracy on 1D and 2D problems, a surrogate model was constructed for the 2D combustion, where five parameters were added to the NN inputs so that the model can learn multi-condition solutions, as shown in Figure 17. The parameterized PINN surrogate model was shown to accurately predict the fields, be applicable to sensitive analysis, and be much more efficient for industrial designs than traditional numerical simulations. Song et al. 308 applied PINNs to 2D laminar premixed flames considering detailed chemistry of CH4, which was accomplished by integrating the flamelet/progress variable (FPV) model. The equations of mass, momentum, $Z$, and $c$ were solved by PINN, during which the thermophysical properties and source term of $c$ were extracted from the pre-tabulated FPV library, as illustrated in Figure 18. The model can accurately solve the fields with and without observation data, demonstrating the feasibility of PINNs to integrate detailed reaction mechanisms using tabulation methods. Using pseudo-time stepping PINNs (PTS-PINNs) 215, Cao et al. 309 solved the premixed and non-premixed flames with a one-step mechanism of CH4 and obtained satisfactory results. Vanilla PINNs failed to solve the premixed combustion with large gradients and tended to predict non-reacting solutions, but PTS-PINNs showed great convergence performance not only for the premixed combustion but also for non-premixed combustion where the thin reaction zones introduced a higher level of complexity.
@@
Figure 17: Schematic of the parameterized combustion field surrogate model based on PINNs. Reprinted from Reference 307, with permission.
@@
Figure 18: Schematic of the PINNs coupled with the FPV model. Reprinted from Reference 308, with permission.
@@
The above studies only involve forward problems or inverse problems of field reconstruction without parameter inference, which is another important aspect of inverse problems. In computational 1D laminar flames, which are important for the study of flame limit, stretch, laminar flame speed ($s_{L}$), and flamelet tabulation, unknown eigenvalues exist even in forward problems, such as $s_{L}$ in freely-propagating premixed flames 4, 7, 8. Wu et al. 115 established the FlamePINN-1D framework to solve both forward and inverse problems of 1D laminar flames in a unified manner, as illustrated in Figure 19. Three cases were solved: freely-propagating premixed flames with simplified and detailed physical models, and counterflow premixed flames with detailed physical models. The detailed physical models are the same as those in Section (3.1) to (3.5) with a one-step global mechanism of CH4 being considered, except that the governing equations are for 1D reacting stagnation flows. A warmup pretraining strategy and some hard physical constraints were proposed. For example, the predicted species mass fractions are transformed by the Softmax function:
@@
where $Y_{k}^{\#}$ is the output of NN. The transformation ensures that the range constraint ($0\leqslant Y_{k}\leqslant 1$) and species conservation ($\sum_{k}Y_{k}=1$, Equation (3.1)) can be exactly satisfied, so it can be used for various combustion problems. $Y^{\#}$ can be viewed as the logarithmic normalization of $Y$ 271, which can mitigate the multiscale of their magnitudes. The results show that for forward problems, FlamePINN-1D can accurately solve the flame fields and simultaneously infer the $s_{L}$ of freely-propagating premixed flames. For inverse problems, FlamePINN-1D can accurately reconstruct the continuous fields and infer the transport properties and chemical kinetics parameters from noisy sparse observations, which indicates the possibility of optimizing chemical mechanisms from observations of laboratory 1D flames. In fact, in their inverse problems where the activation energy was unknown, the chemical kinetics part can be considered as a CRNN with other parameters are known. It is proven that the model tends to get unburned solutions without warmup pretraining, the similar situation to Reference 309. Other adopted strategies, such as hard constraints, thin-layer normalization 88, and modified MLP with random weight factorization (MMLP-RWF) 82, have also been proven to be essential for the model performance.
Figure 19: FlamePINN-1D for solving forward and inverse problems of 1D laminar flames. (a) Schematic of the FlamePINN-1D framework. (b) Case 1: freely-propagating premixed flames with simplified physical models. (c) Case 2: freely-propagating premixed flames with detailed physical models. (d) Case 3: counterflow premixed flames with detailed physical models. Reprinted from Reference 115, with permission.
In addition to the classical PINNs, the PINNs in a broad sense (Table 1) can also be applied to laminar combustion problems. For example, Corrochano et al. 326 developed a physics-aware ROM for an axisymmetric laminar CH4-O2 flame. The simulation data were preprocessed to obtain the proper orthogonal decomposition (POD) modes, then deep learning models were used as a time integrator to advance temporal modes in time, i.e., the models took the AR paradigm instead of the NAR paradigm adopted in classical PINNs (see Section 2.3.1). A physics-aware loss function was introduced which constrained the predicted species mass to be equal to the true ones in each time instant. This loss differed from the classical PINN losses in that it was still a data-driven loss without PDE information, so it was called "physics-aware" rather than "physics-informed", but can still be called "physics-informed" in a broad sense in our extended definition. Two models, including the LSTM and 1D CNN, were tested, demonstrating that the physics-aware loss could slightly improve the model performance.